OpenStudy (anonymous):

Work out and give your answer in standard form, (5 x 10^-8) * (2 x 10^3)?

OpenStudy (anonymous):

5*2*10^(-8+3) = 10^(-4) = 1/10000 = 0.0001

OpenStudy (anonymous):

Ah, I had used a different method in class to work this out, but your method works too- thank you so much, just had to check hehe :)!

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

a^(m) * a(^n) = a^(m+n)

OpenStudy (anonymous):

Ah yes- I remember that!- I had forgotten that formula for about a month haha :P

OpenStudy (anonymous):

np :)